OpenStudy (anonymous):

Two less than 6 times a number is the same as 8 times the number find the number

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Let the required number be x Therefore 6x-2=8x -2= 8x-6x -2= 2x x= -1 hence the required number is -1

OpenStudy (anonymous):

thank you can u help with another?

OpenStudy (anonymous):

Sure,,

OpenStudy (anonymous):

p-21%=$54.60

OpenStudy (anonymous):

p-21%=$54.60 i.e. p-21/100= 54.60 i.e. (100p -21)/100 = 54.60 i.e. 100p-21= 54.60 * 100 i.e. 100p-21= 5460 i.e. 100p= 5460 +21 i.e. 100p= 5481 i.e. p= 5481/100 i.e. p= 54.81

OpenStudy (anonymous):

thanks same answer i got :)

OpenStudy (anonymous):

Very good..